# An All-Atom Generative Model for Designing Protein Complexes

### v2

@@ 3.3.3 Training of Refine Module
The Refine Module $p_{\text{Refine}}(\tilde{\mathbf{S}}_{1},\tilde{\mathbf{T}}_{1}|\hat{\mathbf{S}}_{1},\hat{\mathbf{T}}_{1},t_{S},t_{T},\hat{\mathbf{\chi}})$ is tasked to predict the real protein based on the generated one, $(\hat{\mathbf{S}}_{1},\hat{\mathbf{T}}_{1})$, with the all-atom level information formed with the predicted sidechain torsions $\hat{\mathbf{\chi}}$. Thus, we define a correction loss on the corrected protein $(\tilde{\mathbf{S}}_{1},\tilde{\mathbf{T}}_{1})$ as the learning objective as:
@@ 4.3.2 Multi-Chain Protein Generation
$\Delta\text{G}$downarrow $<$0 $\uparrow$ pLDDT $\uparrow$ ipTM $\uparrow$ Success $\uparrow$ DockQ $\uparrow$ $\geq$0.8 $\uparrow$
@@
$\textrm{APM}_{\text{SFT}}$19.90 69.34 60.36 0.66 29.22% 0.40 11.29
$\textrm{APM}_{\text{zero-shot}}$23.71 62.18 60.97 0.62 27.20% 0.24 0.12
@@
[…]
@@ 4.4.2 Peptide Design
We compare our model with: PepGLAD (Kong et al., 2024), PPFlow (Lin et al., 2024), and DiffPP (Lin et al., 2024). We also include RFDiffusion (Watson et al., 2023), which utilizes ProteinMPNN for sequence design. The peptide candidates are comprehensively evaluated across three key aspects: Functionality, Foldability, and Accuracy. For functionality, we evaluate the binding energy ($\Delta\text{G}$) and the proportion of candidates with $\Delta\text{G}$ below zero, % $<$0. For foldability, we use Boltz-1(wMSA) to fold sequences of generated peptides, then evaluate the folded structure in two confidence metrics: predicted Local Distance Difference Test (pLDDT) and interface predicted Template Modeling (ipTM(Zhang & Skolnick, 2004; Xu & Zhang, 2010) score. We also report a comprehensive metric, defined as the proportion of candidates with both a pLDDT score $\geq 70$ and an ipTM score $\geq 0.8$, donated as Success. For accuracy, we evaluate the DockQ (Basu & Wallner, 2016; Mirabello & Wallner, 2024) score and the proportion of candidates achieving a DockQ score of at least 0.8 ($\geq$0.8), which is the threshold considered as high-quality.
@@ A.2 Discrete Flow Matching
where $\delta\{a,b\}=1$ if $a=b$ and 00 otherwise. This interpolant linearly mixes the clean sequence and the mask state over time.
@@ B.1 Loss for Seq&BB Module
[…]
@@
[…]
@@
[…]
@@ B.4 Training
In training phase I, the Seq&BB Module was trained on 64$\times$×H100 GPUs with 257,000 steps, with a learning rate of 1e-4. The Sidechain Module was trained on 8$\times$×H100 GPUs, accumulating a total of 836,901 steps, also with a learning rate of 1e-4.
@@
In training phase II, APM was trained on 64$\times$×H100 GPUs with 235,000 steps . The learning rate for the Seq&BB Module is set to 1e-5.
@@
In the SFT phase, we used 8$\times$×H100 GPUs to fine-tune APM for antibody design and peptide design. The SFT phase lasted for 1200 epochs for every task. The learning rate for each module is set to 5e-5 and the training cycle is adjusted to 10-1-1.
@@ D.2 Statistical Validation on Folding
We conducted folding with APM and ESM3 using 20 random seeds. For RMSD, ESM3 shows a marginally better mean (4.708±0.094 vs 4.828±0.077) with statistical significance (p $<$ 0.05). For TM-score, APM achieves better performance (0.856±0.002 vs 0.828±0.002) with statistical significance (p $<$ 0.05). The detailed results are shown below with the format of (average±std).
@@ D.3 Multi-Chain Protein
Theoretically, APM supports the generation of protein complexes composed of any number of chains. However, we observed that APM’s performance declines when the generated complex contains more than two chains. This is evident in the increased presence of unstructured backbones or abnormally high ratios of $\alpha$-helix/$\beta$-sheet structures (over 90% of single secondary structure) in the generated complexes. Interestingly, these structures still exhibit no obvious clashes in the binding interface. We do not apply detailed metric calculations since $\Delta\text{G}$ only computes the binding strength between two components. Instead, we present some cases for proteins composed of more chains in Figure 7.
@@ Methods.
Table 10: Performance comparison of antibody design methods on RAbD benchmark. ($\uparrow$/$\downarrow$) indicate whether higher or lower values are better. The best results are shown in bold.
[…]
@@
  • Foldability. We only fold generated sequences with lengths greater than or equal to 10 residues (46 peptide-receptor pairs in total). As indicated in the documentation of Boltz-1 and AlphaFold3, the ipTM scores may not be reliable for sequences that are too short. To ensure reliability, we only consider peptide-receptor pairs where the ground-truth samples achieve successful confidence scores in Boltz-1 (pLDDT70absent70$\geq$ 70≥ 70 and ipTM0.8absent0.8$\geq$ 0.8≥ 0.8.), filtering out cases where even the ground truth structures fail to meet the confidence threshold. The two threshold values are adopted from the official output documentation [^4] https://github.com/google-deepmind/alphafold3/blob/main/docs/output.md of AlphaFold3 and paper (Abramson et al., 2024). Considering the computational cost and MSA retrieval time, we select the top 16 sequences ranked by $\Delta\text{G}$ and perform folding 8 times for each peptide-receptor pair. For each sequence, we select the best folding result, then average these results across all 16 sequences to obtain the final confidence score.
@@ Appendix F Binder Design
Settings. We further explore APM’s zero-shot capabilities in binder design, focusing specifically on longer protein binders rather than short peptides. Following previous works (Watson et al., 2023; Zambaldi et al., 2024), we selected several important targets: Interleukin-7 Receptor-$\alpha$(IL-7RA), SARS-CoV-2 spike protein receptor-binding domain (SC2RBD), MDM2, Programmed Death-1 (PD1), Programmed Death-Ligand 1 (PD-L1), and CD3-epsilon (CD3E). We extracted the corresponding target chains and reference binders from PDB: 3di3, 6m0j, 1ycr, 4zqk, 4z18, and 1xiw, respectively. The evaluation settings remain consistent with peptide design experiments. However, due to uncertainty about which amino acids should be defined as hot-spot residues for these targets, we did not define hot-spot residues for RFDiffusion nor extract pockets for APM. Consequently, we do not report DockQ for this task. Additionally, due to computational resource constraints, we folded only the top 8 sequences rather than 16 as in previous experiments. The results of average metrics are presented in Table 11, with the ’Success’ representing the proportion of samples (out of 8) that achieve both pLDDT $>$ 80 and ipTM $>$ 0.8. We present the generated binders across different methods and their corresponding folded structures with the highest pLDDT sequences in Figure 13.
@@
Figure 13: Visualization of binder design results across six protein targets. From left to right: (1) GroundTruth: native complex structures, (2) APM: structures generated by APM, (3) RFDiffusion: structures generated by RFDiffusion, (4) APM(folded): highest pLDDT APM sequences folded with Boltz, (5) RFDiffusion(folded): highest pLDDT sequences folded with Boltz. The blue regions represent the given targets, while the pink regions show the generated binders. For folded structures, targets are rendered transparent to highlight binders, with binders colored according to AlphaFold’s pLDDT scheme. Each row represents a different PDB target: 3di3, 6m0j, 1ycr, 4zqk, 4z18, and 1xiw.
@@
Discussion. Overall, APM is comparable to RFDiffusion. Although these metrics have been proven by many studies to be predictive of wet lab experimental results (Zambaldi et al., 2024; Bennett et al., 2025), the actual effectiveness still requires validation through wet lab experiments. We observed that both APM and RFdiffusion encounter cases where some samples exhibit lower pLDDT and ipTM scores. The pLDDT score can vary significantly along a protein chain. This means the folding model can be very confident in the structure of some regions of the protein, but less confident in other regions. We hypothesize that the low pLDDT scores stem from the complexity of long binders. Specifically, certain regions may be naturally highly flexible or intrinsically disordered, leading the folding model to assign low pLDDT scores to these residues (as indicated in  (Guo et al., 2022)). Regarding ipTM, we speculate that the lower scores may result from the larger binding interfaces typical of long binders, which often involve multiple contact points or complex features such as convex or polar epitopes, or hydrophobic regions (Zambaldi et al., 2024). These structural complexities and biological properties can contribute to lower ipTM scores.
@@
Target Method $\Delta\text{G}$downarrow $<$0 $\uparrow$ pLDDT ipTM Success
3di3 GroundTruth -23.79 - 95.26 0.85 -
RFDiffusion -50.49 82.50% 87.83 0.30 0%
$\text{APM}_{\text{MPNN}}$80.09 92.50% 83.39 0.29 12.5%
APM -80.10 95.00% 78.91 0.38 12.5%
6m0j GroundTruth -20.11 - 81.55 0.15 -
RFDiffusion -56.10 67.50% 70.90 0.45 0%
$\text{APM}_{\text{MPNN}}$88.99 65.00% 67.82 0.40 12.5%
APM -96.47 67.50% 69.50 0.48 12.5%
@@
$\text{APM}_{\text{MPNN}}$33.27 90.00% 71.10 0.70 50%
@@
$\text{APM}_{\text{MPNN}}$43.33 77.50% 79.10 0.36 0%
@@
4z18 GroundTruth -40.89 - 92.08 0.76 -
RFDiffusion -18.69 57.50% 67.39 0.30 0%
$\text{APM}_{\text{MPNN}}$63.37 55.00% 74.28 0.34 0%
APM -54.24 55.00% 69.28 0.35 0%
@@
$\text{APM}_{\text{MPNN}}$46.96 82.50% 72.08 0.70 12.5%

### v3

@@ 3.3.3 Training of Refine Module
[…]
@@ 4.3.2 Multi-Chain Protein Generation
$\Delta\text{G}$ $\downarrow$ <<0 $\uparrow$ pLDDT $\uparrow$ ipTM $\uparrow$ Success $\uparrow$ DockQ $\uparrow$ $\geq$0.8 $\uparrow$
@@
$\textrm{APM}_{\text{SFT}}$ -19.90 69.34 60.36 0.66 29.22% 0.40 11.29
$\textrm{APM}_{\text{zero-shot}}$ -23.71 62.18 60.97 0.62 27.20% 0.24 0.12
@@
Method AAR (%) $\uparrow$ RMSD $\downarrow$ $\text{E}\downarrow$ $\Delta\text{G}\downarrow$
@@ 4.4.2 Peptide Design
We compare our model with: PepGLAD (Kong et al., 2024), PPFlow (Lin et al., 2024), and DiffPP (Lin et al., 2024). We also include RFDiffusion (Watson et al., 2023), which utilizes ProteinMPNN for sequence design. The peptide candidates are comprehensively evaluated across three key aspects: Functionality, Foldability, and Accuracy. For functionality, we evaluate the binding energy ($\Delta\text{G}$) and the proportion of candidates with $\Delta\text{G}$ below zero, % <<0. For foldability, we use Boltz-1(wMSA) to fold sequences of generated peptides, then evaluate the folded structure in two confidence metrics: predicted Local Distance Difference Test (pLDDT) and interface predicted Template Modeling (ipTM(Zhang & Skolnick, 2004; Xu & Zhang, 2010) score. We also report a comprehensive metric, defined as the proportion of candidates with both a pLDDT score $\geq 70$ and an ipTM score $\geq 0.8$, donated as Success. For accuracy, we evaluate the DockQ (Basu & Wallner, 2016; Mirabello & Wallner, 2024) score and the proportion of candidates achieving a DockQ score of at least 0.8 (% $\geq$0.8), which is the threshold considered as high-quality.
@@ A.2 Discrete Flow Matching
where $\delta\{a,b\}=1$ if $a=b$ and 0 otherwise. This interpolant linearly mixes the clean sequence and the mask state over time.
@@ B.1 Loss for Seq&BB Module
$\mathcal{L}_{\text{consistency\_S}}=\text{KL}\Bigl{(}\log\bigl{(}\text{softmax}(\hat{L}_{1})\bigr{)},\log\bigl{(}\text{softmax}(\hat{L}_{1}^{\prime})\text{.detach()}\bigr{)}\Bigr{)}$ (17)
@@
$\mathcal{L}_{\text{consistency\_T}}=\text{MSE}\Bigl{(}\text{trans}(\hat{\mathbf{T}}_{1}),\text{trans}(\hat{\mathbf{T}}_{1}^{\prime})\text{.detach()}\Bigr{)}+\text{MSE}\Bigl{(}\text{Mat2Vec}\bigl{(}\text{rot}(\hat{\mathbf{T}}_{1})\bigr{)},\text{Mat2Vec}\bigl{(}\text{rot}(\hat{\mathbf{T}}_{1}^{\prime})\bigr{)}\text{.detach()}\Bigr{)}$ (18)
@@
$\mathcal{L}_{\text{consistency}}=t^{2}_{S}\times(\sqrt[2]{{\mathcal{L}_{\text{consistency}\_S}}^{2}+c_{S}^{2}}-c_{S})+t^{2}_{T}\times(\sqrt[2]{{\mathcal{L}_{\text{consistency}\_T}}^{2}+c_{T}^{2}}-c_{T})$ (19)
$c_{S}=0.00054*\sqrt[2]{\text{dim}_{S}},\quad c_{T}=0.00054*\sqrt[2]{\text{dim}_{T}}$ (20)
@@ B.4 Training
[…]
@@
[…]
@@
[…]
@@ D.2 Statistical Validation on Folding
We conducted folding with APM and ESM3 using 20 random seeds. For RMSD, ESM3 shows a marginally better mean (4.708±0.094 vs 4.828±0.077) with statistical significance (p << 0.05). For TM-score, APM achieves better performance (0.856±0.002 vs 0.828±0.002) with statistical significance (p << 0.05). The detailed results are shown below with the format of (average±std).
@@ D.3 Multi-Chain Protein
[…]
@@ Methods.
Table 10: Performance comparison of antibody design methods on RAbD benchmark. ($\uparrow$$\downarrow$) indicate whether higher or lower values are better. The best results are shown in bold.
Method $\text{E}\downarrow$ $\Delta\text{G}\downarrow$
@@
[…]
@@ Appendix F Binder Design
Settings. We further explore APM’s zero-shot capabilities in binder design, focusing specifically on longer protein binders rather than short peptides. Following previous works (Watson et al., 2023; Zambaldi et al., 2024), we selected several important targets: Interleukin-7 Receptor-$\alpha$(IL-7RA), SARS-CoV-2 spike protein receptor-binding domain (SC2RBD), MDM2, Programmed Death-1 (PD1), Programmed Death-Ligand 1 (PD-L1), CD3-epsilon (CD3E), and TNF-$\alpha$. We extracted the corresponding target chains and reference binders from PDB: 7opb, 7zsd, 1ycr, 4zqk, 8znl, 1xiw, and 1tnf respectively. The evaluation settings remain consistent with peptide design experiments. For the TNF-$\alpha$ target, the binder length was set to 50 residues, since no binding PDB structure was found. However, due to uncertainty about which amino acids should be defined as hot-spot residues for these targets, we did not define hot-spot residues for RFDiffusion nor extract pockets for APM. Consequently, we do not report DockQ for this task. Additionally, due to computational resource constraints, we folded only the top 8 sequences rather than 16 as in previous experiments. The results of average metrics are presented in Table 11, with the ’Success’ representing the proportion of samples (out of 8) that achieve both pLDDT >> 80 and ipTM >> 0.8. We present the generated binders across different methods and their corresponding folded structures with the highest pLDDT sequences in Figure 13.
@@
Figure 13: Visualization of binder design results across six protein targets. From left to right: (1) GroundTruth: native complex structures, (2) APM: structures generated by APM, (3) RFDiffusion: structures generated by RFDiffusion, (4) APM(folded): highest pLDDT APM sequences folded with Boltz, (5) RFDiffusion(folded): highest pLDDT sequences folded with Boltz. The blue regions represent the given targets, while the pink regions show the generated binders. For folded structures, targets are rendered transparent to highlight binders, with binders colored according to AlphaFold’s pLDDT scheme. Each row represents a different PDB target: 7opb, 7zsd, 1ycr, 4zqk, 8znl, 1xiw and 1tnf.
@@
Discussion. Overall, APM is comparable to RFDiffusion, except in the case of the trimer target TNF-$\alpha$. Although these metrics have been proven by many studies to be predictive of wet lab experimental results (Zambaldi et al., 2024; Bennett et al., 2025), the actual effectiveness still requires validation through wet lab experiments. We observed that both APM and RFdiffusion encounter cases where some samples exhibit lower pLDDT and ipTM scores. The pLDDT score can vary significantly along a protein chain. This means the folding model can be very confident in the structure of some regions of the protein, but less confident in other regions. We hypothesize that the low pLDDT scores stem from the complexity of long binders. Specifically, certain regions may be naturally highly flexible or intrinsically disordered, leading the folding model to assign low pLDDT scores to these residues (as indicated in  (Guo et al., 2022)). Regarding ipTM, we speculate that the lower scores may result from the larger binding interfaces typical of long binders, which often involve multiple contact points or complex features such as convex or polar epitopes, or hydrophobic regions (Zambaldi et al., 2024). These structural complexities and biological properties can contribute to lower ipTM scores.
@@
Target Method $\Delta\text{G}$ $\downarrow$ <<0 $\uparrow$ pLDDT ipTM Success
7opb GroundTruth -50.06 - 95.58 0.94 -
RFDiffusion -44.23 100% 71.14 0.55 25.00%
$\text{APM}_{\text{MPNN}}$ -50.25 100% 76.60 0.61 12.50%
APM -55.72 100% 73.73 0.46 12.50%
7zsd GroundTruth -36.65 - 81.26 0.25 -
RFDiffusion -43.99 100% 68.33 0.24 0%
$\text{APM}_{\text{MPNN}}$ -63.01 92.50% 62.34 0.38 0%
APM -48.10 92.50% 71.33 0.39 12.50%
@@
$\text{APM}_{\text{MPNN}}$ -33.27 90.00% 71.10 0.70 50%
@@
$\text{APM}_{\text{MPNN}}$ -43.33 77.50% 79.10 0.36 0%
@@
8znl GroundTruth -44.24 - 95.54 0.95 -
RFDiffusion -30.20 97.50% 78.93 0.56 12.50%
$\text{APM}_{\text{MPNN}}$ -52.85 92.50% 78.00 0.59 12.50%
APM -44.14 92.50% 72.11 0.47 0%
@@
$\text{APM}_{\text{MPNN}}$ -46.96 82.50% 72.08 0.70 12.5%